OpenStudy (anonymous):

The base of a ladder should be placed 2 feet from the wall for every 5 feet of ladder length. How can a 15-foot ladder safely reach? Round to the nearest hundreth.

OpenStudy (anonymous):

Because you get 2 feet from wall per 5 feet of ladder, 15 * (2/5) = 3 * 2 = 6 feet from the wall. Assuming you're looking for how high, x^2 + 36 = 225 x = 16.15 feet

OpenStudy (anonymous):

i think you messed up your math. Its 13.75. but thank you for the equation! it helped!

OpenStudy (anonymous):

Oops! My bad, the answer is just \[\sqrt{189}\] which is around 13.75 (: